Compute the mean of the population(ų)2,5,6,7 and 8.


Compute the variance of the population (ò).

1
Expert's answer
2022-03-20T14:38:07-0400

The mean of the population

"\mu=E(X)=(2+5+6+7+8)/5=5.6"

The mean of "X^2":

"E(X^2)=(2^2+5^2+6^2+7^2+8^2)\/5=35.6"

The variance of the population

"\sigma^2=Var(X)=E(X^2)-\mu^2=35.6-5.6^2=4.24"